# LanPaint: Training-Free Diffusion Inpainting with Asymptotically Exact and Fast Conditional Sampling

### v2

@@ E.2 Comparison with BiG score Drift Term
$r_{t}$ $=\mathbb{E}_{p(\mathbf{x}_{t},\mathbf{y}_{t})}\left\|\frac{\mathbf{y}_{t}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}\right\|_{2},$
@@
$\mathbf{s}^{*}_{\mathbf{y}}=(1+\lambda)\,\frac{s_{t}}{r_{t}}\,\mathbf{s}_{\text{cond}}+\mathbf{g}_{\lambda}(\mathbf{x}_{t},\mathbf{y}_{t},t).$
@@
$\mathbf{s}^{*}(\mathbf{x},\mathbf{y},t)=\Big{(}\mathbf{s}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t),\,\mathbf{g}_{\lambda}(\mathbf{x},\mathbf{y},t)\Big{)}+\frac{s_{t}}{r_{t}}\big{(}\mathbf{0},\,\mathbf{s}_{\text{cond}}\big{)}.$
E.3 Perturbation Between Ideal and BiG score Scores
@@
$r(t)$ $=\mathbb{E}_{p_{t}(\mathbf{x},\mathbf{y})}\left\|\frac{\mathbf{y}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}\right\|_{2},$
@@
From the relationship between the score function and the noise term in diffusion models (see equation 39), we have:
$\bm{\epsilon}(\mathbf{x},t)=-\sqrt{1-\bar{\alpha}_{t}}\,\mathbf{s}(\mathbf{x},t),$
@@
$\mathbf{s}(\mathbf{x},t)=-\frac{\bm{\epsilon}(\mathbf{x},t)}{\sqrt{1-\bar{\alpha}_{t}}}.$
@@
$\mathbb{E}_{p(\mathbf{x},t)}\left\|\bm{\epsilon}(\mathbf{x},t)\right\|_{2}=C.$
Under Assumption E.1, we can derive the scaling behavior of $s_{t}$ and $r_{t}$:
@@
$s(t)=\mathbb{E}_{p_{t}(\mathbf{x},\mathbf{y})}\left\|\nabla_{\mathbf{y}}\log p_{t}(\mathbf{x}\mid\mathbf{y})\right\|_{2}\sim\frac{C}{\sqrt{1-\bar{\alpha}_{t}}}.$
@@
$r(t)=\mathbb{E}_{p_{t}(\mathbf{x},\mathbf{y})}\left\|\frac{\mathbf{y}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}\right\|_{2}\sim\frac{C^{\prime}}{1-\bar{\alpha}_{t}},$
@@ 3. Ratio $\frac{s_{t}}{r_{t}}$: Combining the scaling behaviors of $s_{t}$ and $r_{t}$, we obtain:
$\frac{s_{t}}{r_{t}}\sim\frac{C/\sqrt{1-\bar{\alpha}_{t}}}{C^{\prime}/(1-\bar{\alpha}_{t})}=\frac{C}{C^{\prime}}\sqrt{1-\bar{\alpha}_{t}}.$
@@
Now we have shown that the idealized Langevin dynamics equation 57 and the BiG score Langevin dynamics equation 56 differ only by an $\mathcal{O}(\sqrt{1-\bar{\alpha}_{t}})$ perturbation.
E.4 Fokker-Planck Equation Analysis
@@
The Fokker–Planck equation equation 105 describes the time evolution of the probability density $\rho(\mathbf{z},\tau)$ associated with a stochastic process governed by a stochastic differential equation (SDE). For a general SDE of the form:
$dz_{i}=h_{i}(\mathbf{z})\,d\tau+\gamma_{ij}(\mathbf{z})\,dW_{j},$
@@
$\frac{\partial\rho(\mathbf{z},\tau)}{\partial\tau}=\mathcal{L}\rho(\mathbf{z},\tau),$
@@
$\mathcal{L}=-\nabla\cdot\bigl{[}\mathbf{h}(\mathbf{z})\,\cdot\bigr{]}+\frac{1}{2}\nabla\cdot\bigl{[}\mathbf{D}(\mathbf{z})\nabla\cdot\bigr{]}.$
[…]
E.4.1 Fokker-Planck Operator for BiG score
The BiG score dynamics are governed by the SDEs equation 56. The corresponding Fokker-Planck operator for the BiG score is:
$\mathcal{L}_{\text{BiG score}}=-\nabla_{\mathbf{x}}\cdot\bigl{[}\mathbf{s}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t)\,\cdot\bigr{]}-\nabla_{\mathbf{y}}\cdot\bigl{[}\mathbf{g}_{\lambda}(\mathbf{x},\mathbf{y},t)\,\cdot\bigr{]}+\nabla^{2},$
@@ E.4.1 Fokker-Planck Operator for BiG score
The corresponding Fokker-Planck operator for the idealized SDE equation 57 is:
$\mathcal{L}_{\text{ideal}}=-\nabla_{\mathbf{x}_{t}}\cdot\bigl{[}\mathbf{s}_{\mathbf{x}}(\mathbf{x}_{t},\mathbf{y}_{t},t)\,\cdot\bigr{]}-\nabla_{\mathbf{y}_{t}}\cdot\bigl{[}\mathbf{s}^{*}_{\mathbf{y}}(\mathbf{x}_{t},\mathbf{y}_{t},t)\,\cdot\bigr{]}+\nabla^{2}.$
E.4.2 Deviation Between BiG score and Idealized SDE
@@
$\mathbf{s}^{*}_{\mathbf{y}}=(1+\lambda)\,\frac{s_{t}}{r_{t}}\,\mathbf{s}_{\text{cond}}+\mathbf{g}_{\lambda},$
@@
$\mathcal{L}_{\text{BiG score}}-\mathcal{L}_{\text{ideal}}=-\nabla_{\mathbf{y}_{t}}\cdot\left[(1+\lambda)\,\frac{s_{t}}{r_{t}}\,\mathbf{s}_{\text{cond}}\,\cdot\right].$
@@
E.4.3 Analysis Using Dyson’s Formula
To analyze the deviation of the solutions to the Fokker-Planck equations, we use Dyson’s Formula Evans & Morriss (2008), a powerful tool in the study of perturbed differential equations. Dyson’s Formula expresses the solution to a perturbed differential equation in terms of the unperturbed solution. Specifically, for a differential equation of the form
$\frac{\partial\rho(\mathbf{z},\tau)}{\partial\tau}=(\mathcal{L}_{0}+\mathcal{L}_{1})\rho(\mathbf{z},\tau),$
@@ E.4.3 Analysis Using Dyson’s Formula
[…]
This formula allows us to express the solution to the perturbed equation as the sum of the unperturbed solution and a correction term due to the perturbation.
In our case, the Fokker-Planck equation for the BiG score can be viewed as a perturbation of the idealized Fokker-Planck equation. Let $\rho_{\text{ideal}}(\mathbf{x}_{t},\mathbf{y}_{t},\tau)$ be the solution to the idealized Fokker-Planck equation:
$\frac{\partial\rho_{\text{ideal}}}{\partial\tau}=\mathcal{L}_{\text{ideal}}\rho_{\text{ideal}}.$
@@
[…]
@@
[…]
From the analysis in Section E.3, we know that $\frac{s_{t}}{r_{t}}\sim\sqrt{1-\bar{\alpha}_{t}}$. Thus, the deviation term is proportional to $\sqrt{1-\bar{\alpha}_{t}}$.
E.5 Conclusion
@@
[…]
@@ E.6 Supp: P is bounded
[…]
@@
    Spectral Gap: Assumption 2 implies that $\mathcal{L}_{\text{ideal}}$ has a spectral gap $L>0$, or equivalently, $-L$ is the largest non-zero eigenvalue.
@@
    Exponential Damping: The term $e^{\mathcal{L}_{\text{ideal}}(\tau-s)}\nabla_{\mathbf{y}_{t}}\cdot\mathbf{S}$ represents the evolution of the initial condition $\nabla_{\mathbf{y}_{t}}\cdot\mathbf{S}$ from time 00 to $\tau-s$. Suppose
@@
[…]
@@
$|P|\leq\frac{\text{constant}}{L}.$
Appendix F Fast Langevin Dynamics (FLD) with Momentum
@@ A Naive Solver of FLD
Solving this exactly over time interval $[0,\Delta\tau]$ with constant $\mathbf{s}$ yields:
@@
We propose a solution to address the white noise issue: by engineering the FLD dynamics such that for large $\Delta\tau$, $\mathbf{x}$ asymptotically converges to the form specified in equation 68. In contrast to equation 67, which introduces infinitely large noise to $\mathbf{x}$ as $\Delta\tau\to\infty$, the asymptotic behavior for large $\Delta\tau$ should satisfy:
[…]
@@
This process has the desired stationary distribution $\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{x}}_{0},1-\bar{\alpha}_{t})$, rigorously satisfying equation 71 and maintaining the correct noise characteristics. Now, let’s design an exact solver for the FLD equation equation 72, treating $\hat{\mathbf{x}}_{0}$ as a constant.
@@
$\mathbb{E}[\bm{\eta}_{\tau}]=0,\quad\mathbb{E}[\bm{\eta}_{\tau}\bm{\eta}_{\tau^{\prime}}^{T}]=\delta(\tau-\tau^{\prime})\,I.$
Equation equation 76 describes a damped harmonic oscillator with noise, whose behavior is governed by the competition between restoring force $A$ and damping $\Gamma$. The key parameter is the discriminant:
@@
The exact solution to the FLD equation equation 72 with initial conditions $\mathbf{x}(0)$ and $\mathbf{q}(0)$ follows a multivariate normal distribution at time $\tau$:
@@
$\bm{\mu}=\begin{pmatrix}\mathbf{x}+\Big{[}\mathbf{q}(0)\zeta_{2}(\Gamma\tau,\Delta)+(\mathbf{C}-A\mathbf{x})(1-\zeta_{1}(\Gamma\tau,\Delta))\Big{]}\tau\\ \mathbf{q}(0)\Big{[}E(\Gamma\tau,\Delta)-A(1-\zeta_{1}(\Gamma\tau,\Delta))\tau\Big{]}+(\mathbf{C}-A\mathbf{x})(1-E(\Gamma\tau,\Delta))\end{pmatrix}$
$\Sigma=D^{2}\begin{pmatrix}\tau\sigma_{22}(\Gamma\tau,\Delta)&\frac{1}{2}[\Gamma\tau\zeta_{2}(\Gamma\tau,\Delta)]^{2}\\ \frac{1}{2}[\Gamma\tau\zeta_{2}(\Gamma\tau,\Delta)]^{2}&\frac{\Gamma}{2}\sigma_{11}(\Gamma\tau,\Delta)\end{pmatrix}$
@@
Sample $\begin{bmatrix}\mathbf{x}_{\tau}\\ \mathbf{q}_{\tau}\end{bmatrix}\sim\mathcal{N}(\bm{\mu},\Sigma)$ according to equation 78;
@@
For the FLD equation equation 72, the coefficients take specific forms:
@@
where $\mathbf{s}(\mathbf{x})$ is the score function. Substituting these into the general solution equation 78 yields an exact analytical solver for the FLD dynamics ($\hat{\mathbf{x}}_{0}$ treated as constant).
@@
where $\mathbf{C}$ is treated as a constant during a single time step. This freedom allow us to do the following modification to equation 80

### v3

@@ F.2 Comparison with BiG score Drift Term
$r_{t}$ $=\mathbb{E}_{p(\mathbf{x}_{t},\mathbf{y}_{t})}\left\|\frac{\mathbf{y}_{t}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}\right\|_{2},$ (66)
@@
$\mathbf{s}^{*}_{\mathbf{y}}=(1+\lambda)\,\frac{s_{t}}{r_{t}}\,\mathbf{s}_{\text{cond}}+\mathbf{g}_{\lambda}(\mathbf{x}_{t},\mathbf{y}_{t},t).$ (67)
@@
$\mathbf{s}^{*}(\mathbf{x},\mathbf{y},t)=\Big(\mathbf{s}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t),\,\mathbf{g}_{\lambda}(\mathbf{x},\mathbf{y},t)\Big)+\frac{s_{t}}{r_{t}}\big(\mathbf{0},\,\mathbf{s}_{\text{cond}}\big).$ (68)
F.3 Perturbation Between Ideal and BiG score Scores
@@
$r(t)$ $=\mathbb{E}_{p_{t}(\mathbf{x},\mathbf{y})}\left\|\frac{\mathbf{y}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}\right\|_{2},$ (69)
@@
From the relationship between the score function and the noise term in diffusion models (see Eq.41), we have:
$\bm{\epsilon}(\mathbf{x},t)=-\sqrt{1-\bar{\alpha}_{t}}\,\mathbf{s}(\mathbf{x},t),$ (70)
@@
$\mathbf{s}(\mathbf{x},t)=-\frac{\bm{\epsilon}(\mathbf{x},t)}{\sqrt{1-\bar{\alpha}_{t}}}.$ (71)
@@
$\mathbb{E}_{p(\mathbf{x},t)}\left\|\bm{\epsilon}(\mathbf{x},t)\right\|_{2}=C.$ (72)
Under Assumption F.1, we can derive the scaling behavior of $s_{t}$ and $r_{t}$:
@@
$s(t)=\mathbb{E}_{p_{t}(\mathbf{x},\mathbf{y})}\left\|\nabla_{\mathbf{y}}\log p_{t}(\mathbf{x}\mid\mathbf{y})\right\|_{2}\sim\frac{C}{\sqrt{1-\bar{\alpha}_{t}}}.$ (73)
@@
$r(t)=\mathbb{E}_{p_{t}(\mathbf{x},\mathbf{y})}\left\|\frac{\mathbf{y}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}\right\|_{2}\sim\frac{C^{\prime}}{1-\bar{\alpha}_{t}},$ (74)
@@ 3. Ratio $\frac{s_{t}}{r_{t}}$: Combining the scaling behaviors of $s_{t}$ and $r_{t}$, we obtain:
$\frac{s_{t}}{r_{t}}\sim\frac{C/\sqrt{1-\bar{\alpha}_{t}}}{C^{\prime}/(1-\bar{\alpha}_{t})}=\frac{C}{C^{\prime}}\sqrt{1-\bar{\alpha}_{t}}.$ (75)
@@
Now we have shown that the idealized Langevin dynamics Eq.63 and the BiG score Langevin dynamics Eq.59 differ only by an $\mathcal{O}(\sqrt{1-\bar{\alpha}_{t}})$ perturbation. We also provide numerical verification of this analysis in Fig.9.
Figure 9: We analyze the average norm ratio between the ideal score $\mathbf{s}^{*}_{\mathbf{y}}$ (Eq.65) and the component discarded by the BiG score, $\mathbf{s}^{*}_{\mathbf{y}}-\mathbf{g}_{\lambda}$, relative to $\bar{\alpha}_{t}$, in the conditional Gaussian case (Section 5.1). We focus on the right part of the image when $\bar{\alpha}_{t}$ approaches 1 (i.e., $t\rightarrow 0$), which determines the distribution of generated clean image. This ratio decreases at the same rate as $\sqrt{1-\bar{\alpha}_{t}}$, indicating that the BiG score $\mathbf{g}_{\lambda}$ closely approximates the ideal score $\mathbf{s}^{*}_{\mathbf{y}}$ with negligible error as $t\rightarrow 0$. This confirms that the error scales as $\mathcal{O}(\sqrt{1-\bar{\alpha}_{t}})$, laying the foundation for Theorem 4.1.
F.4 Fokker-Planck Equation Analysis
@@
The Fokker–Planck equation Eq.147 describes the time evolution of the probability density $\rho(\mathbf{z},\tau)$ associated with a stochastic process governed by a stochastic differential equation (SDE). For a general SDE of the form:
$dz_{i}=h_{i}(\mathbf{z})\,d\tau+\gamma_{ij}(\mathbf{z})\,dW_{j},$ (76)
@@
$\frac{\partial\rho(\mathbf{z},\tau)}{\partial\tau}=\mathcal{L}\rho(\mathbf{z},\tau),$ (77)
@@
$\mathcal{L}=-\nabla\cdot\bigl[\mathbf{h}(\mathbf{z})\,\cdot\bigr]+\frac{1}{2}\nabla\cdot\bigl[\mathbf{D}(\mathbf{z})\nabla\cdot\bigr].$ (78)
Here, $\mathbf{h}(\mathbf{z})$ is the drift vector, $\mathbf{D}(\mathbf{z})=\gamma(\mathbf{z})\gamma(\mathbf{z})^{\top}$ is the diffusion matrix, and $\nabla\cdot$ denotes the divergence operator.
F.4.1 Fokker-Planck Operator for BiG score
The BiG score dynamics are governed by the SDE Eq.59. The corresponding Fokker-Planck operator for the BiG score is:
$\mathcal{L}_{\text{BiG score}}=-\nabla_{\mathbf{x}}\cdot\bigl[\mathbf{s}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t)\,\cdot\bigr]-\nabla_{\mathbf{y}}\cdot\bigl[\mathbf{g}_{\lambda}(\mathbf{x},\mathbf{y},t)\,\cdot\bigr]+\nabla^{2},$ (79)
@@ F.4.1 Fokker-Planck Operator for BiG score
The corresponding Fokker-Planck operator for the idealized SDE Eq.63 is:
$\mathcal{L}_{\text{ideal}}=-\nabla_{\mathbf{x}_{t}}\cdot\bigl[\mathbf{s}_{\mathbf{x}}(\mathbf{x}_{t},\mathbf{y}_{t},t)\,\cdot\bigr]-\nabla_{\mathbf{y}_{t}}\cdot\bigl[\mathbf{s}^{*}_{\mathbf{y}}(\mathbf{x}_{t},\mathbf{y}_{t},t)\,\cdot\bigr]+\nabla^{2}.$ (80)
F.4.2 Deviation Between BiG score and Idealized SDE
@@
$\mathbf{s}^{*}_{\mathbf{y}}=(1+\lambda)\,\frac{s_{t}}{r_{t}}\,\mathbf{s}_{\text{cond}}+\mathbf{g}_{\lambda},$ (81)
@@
$\mathcal{L}_{\text{BiG score}}-\mathcal{L}_{\text{ideal}}=-\nabla_{\mathbf{y}_{t}}\cdot\left[(1+\lambda)\,\frac{s_{t}}{r_{t}}\,\mathbf{s}_{\text{cond}}\,\cdot\right].$ (82)
@@
F.4.3 Analysis Using Dyson’s Formula
To analyze the deviation of the solutions to the Fokker-Planck equations, we use Dyson’s Formula (Evans & Morriss, 2008), a powerful tool in the study of perturbed differential equations. Dyson’s Formula expresses the solution to a perturbed differential equation in terms of the unperturbed solution. Specifically, for a differential equation of the form
$\frac{\partial\rho(\mathbf{z},\tau)}{\partial\tau}=(\mathcal{L}_{0}+\mathcal{L}_{1})\rho(\mathbf{z},\tau),$ (83)
@@ F.4.3 Analysis Using Dyson’s Formula
$\rho(\mathbf{z},\tau)=e^{(\mathcal{L}_{0}+\mathcal{L}_{1})\tau}\rho(\mathbf{z},0)=e^{\mathcal{L}_{0}\tau}\rho(\mathbf{z},0)+\int_{0}^{\tau}e^{\mathcal{L}_{0}(\tau-s)}\mathcal{L}_{1}\rho(\mathbf{z},s)\,ds.$ (84)
This formula allows us to express the solution to the perturbed Eq.83 as the sum of the unperturbed solution and a correction term due to the perturbation.
In our case, the Fokker-Planck equation for the BiG score can be viewed as a perturbation of the Fokker-Planck equation for the idealized SDE. Let $\rho_{\text{ideal}}(\mathbf{x}_{t},\mathbf{y}_{t},\tau)$ be the solution to the idealized Fokker-Planck equation:
$\frac{\partial\rho_{\text{ideal}}}{\partial\tau}=\mathcal{L}_{\text{ideal}}\rho_{\text{ideal}}.$ (85)
@@
$\rho_{\text{BiG score}}(\mathbf{x}_{t},\mathbf{y}_{t},\tau)=\rho_{\text{ideal}}(\mathbf{x}_{t},\mathbf{y}_{t},\tau)+\int_{0}^{\tau}e^{\mathcal{L}_{\text{ideal}}(\tau-s)}\left(\mathcal{L}_{\text{BiG score}}-\mathcal{L}_{\text{ideal}}\right)\rho_{\text{BiG score}}(\mathbf{x}_{t},\mathbf{y}_{t},s)\,ds.$ (86)
@@
$\rho_{\text{BiG score}}(\mathbf{x}_{t},\mathbf{y}_{t},\tau)=\rho_{\text{ideal}}(\mathbf{x}_{t},\mathbf{y}_{t},\tau)-(1+\lambda)\,\frac{s_{t}}{r_{t}}\int_{0}^{\tau}e^{\mathcal{L}_{\text{ideal}}(\tau-s)}\nabla_{\mathbf{y}_{t}}\cdot\left[\mathbf{s}_{\text{cond}}\,\rho_{\text{BiG score}}(\mathbf{x}_{t},\mathbf{y}_{t},s)\right]\,ds.$ (87)
From the analysis in Section F.3, we know that $\frac{s_{t}}{r_{t}}\sim\sqrt{1-\bar{\alpha}_{t}}$. Thus, the deviation term is proportional to $\sqrt{1-\bar{\alpha}_{t}}$.
F.5 Conclusion
@@
$P=\int_{0}^{\tau}e^{\mathcal{L}_{\text{ideal}}(\tau-s)}\nabla_{\mathbf{y}_{t}}\cdot\left[\mathbf{s}_{\text{cond}}\,\rho_{\text{BiG score}}(\mathbf{x}_{t},\mathbf{y}_{t},s)\right]\,ds$ (88)
@@ F.6 Supp: P is bounded
$P=\int_{0}^{\tau}e^{\mathcal{L}_{\text{ideal}}(\tau-s)}\nabla_{\mathbf{y}_{t}}\cdot\left[\mathbf{s}_{\text{cond}}\,\rho_{\text{BiG score}}(\mathbf{x}_{t},\mathbf{y}_{t},s)\right]\,ds$ (89)
@@
    Spectral Gap: Assumption 2 implies that $\mathcal{L}_{\text{ideal}}$ has a spectral gap $L>0$, or equivalently, L-L is the largest non-zero eigenvalue.
@@
    Exponential Damping: The term $e^{\mathcal{L}_{\text{ideal}}(\tau-s)}\nabla_{\mathbf{y}_{t}}\cdot\mathbf{S}$ represents the evolution of the initial condition $\nabla_{\mathbf{y}_{t}}\cdot\mathbf{S}$ from time 0 to $\tau-s$. Suppose
@@
$\int e^{\mathcal{L}_{\text{ideal}}(\tau-s)}\nabla_{\mathbf{y}_{t}}\cdot\mathbf{S}\,d\mathbf{y}_{t}=0$ (93)
@@
$|P|\leq\frac{\text{constant}}{L}.$ (94)
Appendix G Fast Langevin Dynamics (FLD) with Momentum
@@ A Naive Solver of FLD
Solving this exactly over time interval $0,\Delta\tau]$ with constant $\mathbf{s}$ yields:
@@
We propose a solution to address the white noise issue: by engineering the FLD dynamics such that for large $\Delta\tau$, $\mathbf{x}$ asymptotically converges to the form specified in Eq.104. In contrast to Eq.103, which introduces infinitely large noise to $\mathbf{x}$ as $\Delta\tau\to\infty$, the asymptotic behavior for large $\Delta\tau$ should satisfy:
$\mathbf{x}\sim\mathcal{N}\left(\mathbf{x}\mid\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{x}}_{0},\,1-\bar{\alpha}_{t}\right)$ (107)
@@
This process has the desired stationary distribution $\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{x}}_{0},1-\bar{\alpha}_{t})$, rigorously satisfying Eq.107 and maintaining the correct noise characteristics. Now, let’s design an exact solver for the FLD equation Eq.108, treating $\hat{\mathbf{x}}_{0}$ as a constant.
@@
$\mathbb{E}[\bm{\eta}_{\tau}]=0,\quad\mathbb{E}[\bm{\eta}_{\tau}\bm{\eta}_{\tau^{\prime}}^{T}]=\delta(\tau-\tau^{\prime})\,I.$ (113)
Equation Eq.112 describes a damped harmonic oscillator with noise, whose behavior is governed by the competition between restoring force $A$ and damping $\Gamma$. The key parameter is the discriminant:
@@
The exact solution to the FLD equation Eq.108 with initial conditions $\mathbf{x}(0)$ and $\mathbf{q}(0)$ follows a multivariate normal distribution at time $\tau$:
@@
$\bm{\mu}=\begin{pmatrix}\mathbf{x}+\Big[\mathbf{q}(0)\zeta_{2}(\Gamma\tau,\Delta)+(\mathbf{C}-A\mathbf{x})(1-\zeta_{1}(\Gamma\tau,\Delta))\Big]\tau\\ \mathbf{q}(0)\Big[E(\Gamma\tau,\Delta)-A(1-\zeta_{1}(\Gamma\tau,\Delta))\tau\Big]+(\mathbf{C}-A\mathbf{x})(1-E(\Gamma\tau,\Delta))\end{pmatrix}$ (116)
$\Sigma=D^{2}\begin{pmatrix}\tau\sigma_{22}(\Gamma\tau,\Delta)&\frac{1}{2}[\Gamma\tau\zeta_{2}(\Gamma\tau,\Delta)]^{2}\\ \frac{1}{2}[\Gamma\tau\zeta_{2}(\Gamma\tau,\Delta)]^{2}&\frac{\Gamma}{2}\sigma_{11}(\Gamma\tau,\Delta)\end{pmatrix}$ (117)
@@
Sample $\begin{bmatrix}\mathbf{x}_{\tau}\\ \mathbf{q}_{\tau}\end{bmatrix}\sim\mathcal{N}(\bm{\mu},\Sigma)$ according to Eq.115;
@@
For the FLD equation Eq.108, the coefficients take specific forms:
@@
where $\mathbf{s}(\mathbf{x})$ is the score function. Substituting these into the general solution Eq.115 yields an exact analytical solver for the FLD dynamics ($\hat{\mathbf{x}}_{0}$ treated as constant).
@@
where $\mathbf{C}$ is treated as a constant during a single time step. This freedom allow us to do the following modification to Eq.119